solve 4(x+1)= 2(x-3)-(1-2x)
Found 2 solutions by Mathtut, stanbon:
Answer by Mathtut(3670)

4(x+1)=2(x-3)-(1-2x)---->distribute
:
4x+4=2x-6-1+2x---->collect terms on the right side
:
4x+4=4x-7---->which isnt possible!!!!!!

Answer by stanbon(75887)

solve 4(x+1)= 2(x-3)-(1-2x)
4x+4 = 2x-6-1+2x
4x+4 = 4x-7
4 = -7
-----------
That is a contradiction which comes about because we assume
all those x's can be replaced by some number. They can't.
The problem has no solution.
-----------------
Cheers,
Stan H.